A botanist is studying the distribution of daisies in a field. The field is divided into a number of equal sized squares. The mean number of daisies per square is assumed to be 3.

The daisies are distributed randomly throughout the field. Find the probability that, in a randomly chosen square there will be,

a. More than 2 daisies

b. Either 5 or 6 daisies.


1
Expert's answer
2020-12-24T16:49:31-0500

a. The mean no. of diasies are 3. It follows a poisson distribution.

X ~ Poi(λ = 3)

"P(X=x) = \frac{e^λλ ^x}{x!} \\

= \frac{e^{-3}3^x}{x!} \\

P(X>2) = 1 -P(X≤2) \\

= 1 - [P(X=0) + P(X=1) + p(X=2)] \\

= 1 -( \frac{e^{-3}3^0}{0!} + \frac{e^{-3}3^1}{1!} + \frac{e^{-3}3^2}{2!}) \\

= 1 -(0.0498+0.1494+0.224) \\

P(X>2) = 0.5768"

b. Either 5 or 6 daisies

"P(either \; 5 \; or \; 6 \; daisies ) = P(X=5) + P(X=6) \\

= 0.1008 + 0.0504 \\

= 0.1512"